Assume we have the following information about the demand for WIU T-shirts.

Qwiu = 400 - 25Pwiu + 0.002m + 5Pss + 0.02POP

Create a graph showing the simplified demand equation, intercepts, and current market conditions (m=$45,000 Pss=$60, POP=22,00). Determine the quanity demand when price of tshirts = $30
Answer #1

Answer

Qwiu = 400 - 25Pwiu + 0.002m + 5Pss + 0.02POP

m=$45,000 Pss=$60, POP=22,00

Qwiu = 400 - 25Pwiu + 0.002*45000 + 5*60+ 0.02*2200

Qwiu=834-25Pwiu ....................... the simplified demand equation

=============

to graph it we need two points on the curve and then join it.

Pwiu=0

then

Qwiu=834-25*0=834

Qwiu=0 then

0=834-25Pwiu

Pwiu=834/25=33.36

so the two points are

(834,0) and (0,33.36)

using this, the graph is:

40+ price (0, 33.36) 30+ 20+ demand curve 10 (834, 0) 700 100 200 300 400 500 600 800 900 quantity

==========

price of tshirts = $30

then

Qwiu=834-25*30=84

the quantity demanded is 84 units

========

current market condition at P=$30

40+ price (0, 33.36) current market condition at P=$30 30+ (84, 30) 20+ demand 10+ (834, 0) 100 0 200 300 400 500 600 700 800